OpenStudy (anonymous):

I have an obtuse triangle with an angle of 125 degrees, one side is 34.87cm and another 30cm, how do I find the length of the unknown side which is opposite the angle?

OpenStudy (dumbcow):

law of cosines \[c^2 = a^2 +b^2 -2ab \cos C\]

OpenStudy (anonymous):

Ahhhh ok, thanks so very much :)

OpenStudy (anonymous):

Would it be cos of 125?

OpenStudy (dumbcow):

yes

OpenStudy (anonymous):

mmmm I am coming up with a huge number LOL

OpenStudy (anonymous):

Ok got it now, thanks again :)